Problem 3-12 (Algo) Predetermined Overhead Rate; Disposing of Underapplied or Overapplied Overhead (LO3-4] Luzadis Company maRequired 1 Required 2 Required 3 Required 4 Compute the underapplied or overapplied overhead.Journal entry worksheet < 1 Record the entry to close the balance in the manufacturing overhead account to the cost of goodsView transaction list Journal entry worksheet < 1 Record the allocation of the underapplied/overapplied overhead to various aRequired 1 Required 2 Required 3 Required 4 How much higher or lower will net operating income be if the underapplied or over
Answer #1

1) Overhead rate = 560000/70000 = 8 per MH

Applied overhead = 61000*8 = 488000

Actual overhead = 514000

Under applied overhead = 514000-488000 = 26000

2) Journal entry

No General Journal Debit Credit
Cost of goods sold 26000
Manufacturing overhead 26000

3) Journal entry

No General Journal Debit Credit
Work in process 2600
Finished goods (26000*19%) 4940
Cost of goods sold 18460
Manufacturing overhead 26000

4) Net operating income will be higher (2600+4940) = 7540 if the under applied overhead is allocated rather than closed entirely to cost of goods sold